# neurips_data_2024.bib

### v1

@@ Appendix D Detailed AUC-ROC and average precision results
Can. P. 52.8 ± 0.5 61.8 ± 1.1 68.5 ± 2.1 67.7 ± 1.6 63.9 ± 1.3 63.8 ± 0.0 64.2 ± 2.0 77.1 ± 0.4 97.1 ± 0.7
Flights 66.7 ± 3.3 66.8 ± 1.6 68.3 ± 1.8 72.7 ± 0.2 63.9 ± 0.9 70.5 ± 0.0 70.8 ± 0.5 71.2 ± 0.7 68.9 ± 1.8
Cont. 94.2 ± 1.3 95.5 ± 0.3 96.3 ± 1.1 96.0 ± 0.3 84.5 ± 0.2 88.8 ± 0.0 94.6 ± 0.8 94.2 ± 0.1 97.7 ± 0.1
Appendix E Global performance scores
The results presented in Table 3 and Table 4 assign the individual scores of each time window the same weight and then compute the mean over all scores to get the final score. This score measures the model performance across time, i.e. it is equally important for a model to perform well in periods that only have a few edge occurrences as well as in periods where many edges occur. In some scenarios, the focus might not be to forecast the existence of edges in all time windows equally well but instead forecast for all edges equally well. In the following, we investigate the model performance of link forecasting compared to link prediction using this perspective of model performance.
The results are presented in Table 13 for continuous-time temporal graphs and in Table 14 for discrete-time temporal graphs (corresponding average precision in Table 17 and Table 18). In contrast to the results presented in the main part of this work, the scores are computed over all edges instead of per time window and then averaged. As we can see, the changes between link forecasting and link prediction are less expressed if every edge is weighted the same instead of every time window. Nevertheless, we can still observe the patterns discussed above although not as distinct.
Table 13: Test AUC-ROC scores for link forecasting and the relative change compared to link prediction for continuous-time graphs on the same trained models (standard deviations in Table 15). We compute the AUC-ROC score over all edges instead of per time window or batch as in Table 3. The last row/column provides mean $\mu$ and standard deviation $\sigma$ of the absolute relative change per column/row.
Enron 76.8( $\uparrow$1.1%) 73.4( $\uparrow$0.2%) 68.4( $\uparrow$0.2%) 58.9( $\uparrow$0.2%) 66.7( $\uparrow$0.5%) 78.5( $\downarrow$1.6%) 68.0( $\uparrow$0.5%) 81.1( $\uparrow$0.8%) 76.8( $\uparrow$0.5%) 0.6%±0.5%
UCI 85.4( $\uparrow$3.5%) 60.7( $\uparrow$18.1%) 64.0( $\uparrow$1.5%) 59.5( $\downarrow$0.2%) 57.9( $\downarrow$0.5%) 71.3( $\uparrow$3.1%) 59.9( $\downarrow$0.1%) 80.5( $\downarrow$0.1%) 76.4( $\uparrow$0.2%) 3.0%±5.8%
MOOC 83.9( $\downarrow$0.9%) 79.5( $\downarrow$1.3%) 88.1( $\downarrow$0.5%) 82.2( $\uparrow$0.1%) 70.5( $\uparrow$0.2%) 59.9( $\downarrow$3.2%) 72.3( $\uparrow$0.0%) 74.2( $\uparrow$0.1%) 81.0( $\uparrow$0.0%) 0.7%±1.0%
Wiki. 81.6( $\downarrow$0.2%) 78.3( $\downarrow$0.1%) 84.1( $\downarrow$0.0%) 83.4( $\downarrow$0.0%) 71.6( $\uparrow$0.1%) 77.3( $\uparrow$0.3%) 85.1( $\downarrow$0.0%) 87.7( $\downarrow$0.0%) 80.2( $\uparrow$0.2%) 0.1%±0.1%
LastFM 76.6( $\downarrow$0.6%) 70.2( $\downarrow$1.4%) 78.2( $\downarrow$0.3%) 68.5( $\uparrow$0.0%) 68.0( $\downarrow$0.0%) 78.0( $\downarrow$0.2%) 64.3( $\downarrow$0.0%) 66.1( $\downarrow$0.0%) 78.9( $\uparrow$0.0%) 0.3%±0.5%
Myket 64.0( $\uparrow$0.1%) 64.0( $\downarrow$0.0%) 60.7( $\downarrow$0.1%) 57.4( $\downarrow$0.3%) 32.6( $\uparrow$0.3%) 52.0( $\uparrow$0.0%) 58.4( $\downarrow$0.1%) 59.4( $\downarrow$0.1%) 32.9( $\uparrow$0.3%) 0.1%±0.1%
Social 90.4( $\downarrow$0.6%) 91.1( $\downarrow$1.4%) 91.5( $\downarrow$0.1%) 92.7( $\uparrow$0.0%) 87.8( $\uparrow$0.1%) 86.0( $\uparrow$0.2%) 95.3( $\uparrow$0.1%) 94.1( $\uparrow$0.0%) 97.5( $\uparrow$0.0%) 0.3%±0.5%
Reddit 80.5( $\downarrow$0.1%) 79.5( $\downarrow$0.1%) 80.3( $\downarrow$0.1%) 78.6( $\downarrow$0.1%) 80.2( $\downarrow$0.1%) 78.5( $\downarrow$0.2%) 76.2( $\downarrow$0.1%) 77.1( $\downarrow$0.1%) 80.1( $\downarrow$0.0%) 0.1%±0.1%
$\mu\pm\sigma$ 0.9%±1.1% 2.8%±6.2% 0.4%±0.5% 0.1%±0.1% 0.2%±0.2% 1.1%±1.4% 0.1%±0.2% 0.1%±0.3% 0.1%±0.2%
Table 14: Test AUC-ROC scores for discrete-time temporal graphs as in Table 13. All results with standard deviations are listed in Table 16.
UN V. 56.3( $\downarrow$25.5%) 53.3( $\downarrow$28.7%) 52.0( $\downarrow$25.9%) 54.3( $\uparrow$2.8%) 53.8( $\uparrow$7.4%) 89.7( $\uparrow$0.1%) 53.4( $\uparrow$0.7%) 57.1( $\uparrow$1.4%) 63.9( $\uparrow$3.2%) 10.6%±12.3%
US L. 52.5( $\downarrow$7.1%) 61.8( $\downarrow$22.1%) 57.7( $\downarrow$31.2%) 78.6( $\uparrow$0.1%) 82.0( $\uparrow$0.1%) 68.4( $\uparrow$1.3%) 75.4( $\downarrow$0.1%) 90.4( $\uparrow$0.3%) 89.4( $\uparrow$0.2%) 6.9%±11.6%
UN Tr. 57.6( $\downarrow$13.1%) 50.4( $\downarrow$20.3%) 54.4( $\downarrow$14.0%) 64.1( $\uparrow$3.9%) 67.6( $\uparrow$4.6%) 85.6( $\downarrow$1.0%) 63.7( $\uparrow$4.5%) 68.6( $\uparrow$3.4%) 70.7( $\uparrow$3.5%) 7.6%±6.5%
Can. P. 64.0( $\uparrow$0.5%) 64.6( $\downarrow$3.2%) 72.7( $\downarrow$1.4%) 72.3( $\uparrow$0.5%) 68.1( $\uparrow$0.0%) 61.5( $\downarrow$2.6%) 68.3( $\downarrow$0.1%) 81.7( $\uparrow$0.1%) 83.7( $\downarrow$14.3%) 2.5%±4.6%
Flights 67.3( $\downarrow$3.0%) 65.6( $\downarrow$4.7%) 68.1( $\downarrow$1.0%) 72.6( $\downarrow$0.0%) 65.2( $\uparrow$0.3%) 74.6( $\uparrow$0.0%) 70.5( $\downarrow$0.1%) 70.6( $\downarrow$0.1%) 68.5( $\downarrow$0.5%) 1.1%±1.7%
Cont. 93.3( $\downarrow$1.0%) 94.1( $\downarrow$1.4%) 95.6( $\downarrow$0.5%) 95.3( $\downarrow$0.0%) 83.4( $\uparrow$0.1%) 92.2( 0.0%) 94.7( $\uparrow$0.5%) 93.7( $\downarrow$0.0%) 97.2( $\downarrow$0.0%) 0.4%±0.5%
$\mu\pm\sigma$ 8.4%±9.6% 13.4%±11.7% 12.3%±13.6% 1.2%±1.7% 2.1%±3.2% 0.8%±1.0% 1.0%±1.7% 0.9%±1.3% 3.6%±5.5%
Table 15: Test AUC-ROC scores for link forecasting and link prediction averaged over 5 runs with standard deviations on continuous-time temporal graphs.
Eval Dataset JODIE DyRep TGN TGAT CAWN EdgeBank TCL GraphMixer DyGFormer
Forec. Enron 76.8 ± 3.9 73.4 ± 2.7 68.4 ± 3.4 58.9 ± 1.3 66.7 ± 0.5 78.5 ± 0.0 68.0 ± 5.7 81.1 ± 0.8 76.8 ± 0.5
UCI 85.4 ± 1.0 60.7 ± 2.7 64.0 ± 1.1 59.5 ± 1.5 57.9 ± 0.6 71.3 ± 0.0 59.9 ± 0.8 80.5 ± 0.8 76.4 ± 0.5
MOOC 83.9 ± 3.4 79.5 ± 4.1 88.1 ± 2.0 82.2 ± 0.6 70.5 ± 1.2 59.9 ± 0.0 72.3 ± 0.6 74.2 ± 1.4 81.0 ± 9.0
Wiki. 81.6 ± 0.4 78.3 ± 0.4 84.1 ± 0.6 83.4 ± 0.2 71.6 ± 0.8 77.3 ± 0.0 85.1 ± 0.5 87.7 ± 0.3 80.2 ± 1.6
LastFM 76.6 ± 0.5 70.2 ± 1.2 78.2 ± 3.0 68.5 ± 0.8 68.0 ± 0.3 78.0 ± 0.0 64.3 ± 6.0 66.1 ± 1.7 78.9 ± 0.6
Myket 64.0 ± 2.1 64.0 ± 2.7 60.7 ± 2.3 57.4 ± 0.5 32.6 ± 0.4 52.0 ± 0.0 58.4 ± 2.0 59.4 ± 0.4 32.9 ± 1.0
Social 90.4 ± 2.6 91.1 ± 1.0 91.5 ± 3.3 92.7 ± 0.5 87.8 ± 0.1 86.0 ± 0.0 95.3 ± 0.2 94.1 ± 0.2 97.5 ± 0.1
Reddit 80.5 ± 0.2 79.5 ± 0.8 80.3 ± 0.4 78.6 ± 0.7 80.2 ± 0.3 78.5 ± 0.0 76.2 ± 0.4 77.1 ± 0.5 80.1 ± 1.1
Pred. Enron 76.0 ± 3.0 73.2 ± 2.3 68.3 ± 2.9 58.8 ± 1.2 66.4 ± 0.4 79.8 ± 0.0 67.6 ± 5.6 80.5 ± 0.8 76.4 ± 0.5
UCI 82.5 ± 1.3 51.4 ± 7.7 63.0 ± 1.3 59.6 ± 1.5 58.2 ± 0.6 69.1 ± 0.0 60.0 ± 0.9 80.7 ± 0.8 76.2 ± 0.6
MOOC 84.6 ± 3.1 80.5 ± 3.2 88.5 ± 1.6 82.1 ± 0.6 70.3 ± 1.3 61.9 ± 0.0 72.3 ± 0.6 74.1 ± 1.4 81.0 ± 9.0
Wiki. 81.7 ± 0.4 78.4 ± 0.4 84.1 ± 0.6 83.4 ± 0.2 71.5 ± 0.8 77.1 ± 0.0 85.2 ± 0.5 87.8 ± 0.3 80.0 ± 1.6
LastFM 77.1 ± 0.7 71.2 ± 1.1 78.4 ± 2.7 68.5 ± 0.7 68.1 ± 0.3 78.2 ± 0.0 64.3 ± 6.0 66.1 ± 1.7 78.9 ± 0.6
Myket 63.9 ± 2.1 64.0 ± 2.7 60.8 ± 2.3 57.5 ± 0.4 32.5 ± 0.4 51.9 ± 0.0 58.4 ± 2.0 59.5 ± 0.4 32.8 ± 1.0
Social 91.0 ± 2.4 92.4 ± 0.4 91.5 ± 3.5 92.7 ± 0.5 87.7 ± 0.1 85.8 ± 0.0 95.2 ± 0.2 94.1 ± 0.2 97.4 ± 0.1
Reddit 80.6 ± 0.1 79.5 ± 0.8 80.4 ± 0.4 78.7 ± 0.6 80.2 ± 0.3 78.6 ± 0.0 76.2 ± 0.4 77.1 ± 0.5 80.2 ± 1.1
Table 16: Test AUC-ROC scores for link forecasting and link prediction averaged over 5 runs with standard deviations on discrete-time temporal graphs.
Eval Dataset JODIE DyRep TGN TGAT CAWN EdgeBank TCL GraphMixer DyGFormer
Forec. UN V. 56.3 ± 1.4 53.3 ± 0.8 52.0 ± 7.2 54.3 ± 1.4 53.8 ± 2.1 89.7 ± 0.0 53.4 ± 1.1 57.1 ± 1.6 63.9 ± 1.7
UN Tr. 57.6 ± 3.3 50.4 ± 1.2 54.4 ± 1.5 64.1 ± 1.3 67.6 ± 1.2 85.6 ± 0.0 63.7 ± 1.6 68.6 ± 2.6 70.7 ± 2.6
Can. P. 64.0 ± 0.8 64.6 ± 7.5 72.7 ± 2.7 72.3 ± 2.6 68.1 ± 1.0 61.5 ± 0.0 68.3 ± 3.6 81.7 ± 0.9 83.7 ± 3.9
Flights 67.3 ± 2.0 65.6 ± 1.8 68.1 ± 1.7 72.6 ± 0.2 65.2 ± 1.7 74.6 ± 0.0 70.5 ± 0.1 70.6 ± 0.3 68.5 ± 1.3
Cont. 93.3 ± 1.9 94.1 ± 0.5 95.6 ± 0.5 95.3 ± 0.3 83.4 ± 0.1 92.2 ± 0.0 94.7 ± 0.5 93.7 ± 0.1 97.2 ± 0.0
Pred. UN V. 75.6 ± 1.9 74.8 ± 1.2 70.2 ± 5.8 52.8 ± 1.6 50.1 ± 1.6 89.5 ± 0.0 53.0 ± 1.6 56.3 ± 2.0 61.9 ± 1.6
US L. 56.5 ± 1.9 79.3 ± 1.0 84.0 ± 2.2 78.5 ± 7.8 81.9 ± 4.0 67.5 ± 0.0 75.4 ± 5.5 90.2 ± 1.5 89.3 ± 0.9
UN Tr. 66.3 ± 3.0 63.2 ± 1.9 63.2 ± 1.2 61.7 ± 1.3 64.7 ± 1.3 86.4 ± 0.0 60.9 ± 1.3 66.3 ± 2.5 68.3 ± 2.3
Can. P. 63.6 ± 0.7 66.8 ± 2.4 73.7 ± 3.5 72.0 ± 2.6 68.1 ± 1.0 63.1 ± 0.0 68.4 ± 3.6 81.6 ± 1.0 97.7 ± 0.6
Flights 69.4 ± 2.3 68.9 ± 1.0 68.7 ± 1.6 72.6 ± 0.2 65.0 ± 1.4 74.6 ± 0.0 70.6 ± 0.1 70.7 ± 0.3 68.9 ± 1.1
Cont. 94.3 ± 1.2 95.4 ± 0.3 96.1 ± 0.7 95.3 ± 0.3 83.3 ± 0.1 92.2 ± 0.0 94.3 ± 1.0 93.7 ± 0.1 97.3 ± 0.0
Table 17: Average Precision scores computed as in Table 13 for ROC-AUC scores on continuous-time temporal graphs. For a full list of results with standard deviations, see Table 19.
Enron 69.8( $\uparrow$1.1%) 68.2( $\downarrow$0.4%) 65.3( $\uparrow$0.5%) 62.7( $\uparrow$0.1%) 66.8( $\uparrow$0.4%) 75.7( $\downarrow$1.2%) 70.7( $\uparrow$0.9%) 81.5( $\uparrow$1.2%) 77.1( $\uparrow$0.8%) 0.7%±0.4%
UCI 85.1( $\uparrow$6.8%) 56.6( $\uparrow$18.1%) 71.9( $\uparrow$0.4%) 69.0( $\downarrow$0.1%) 65.6( $\uparrow$0.0%) 66.9( $\uparrow$3.1%) 69.6( $\downarrow$0.1%) 85.9( $\downarrow$0.2%) 81.2( $\uparrow$0.2%) 3.2%±6.0%
MOOC 82.8( $\downarrow$0.5%) 76.2( $\downarrow$0.6%) 86.3( $\downarrow$0.7%) 84.4( $\uparrow$0.0%) 73.6( $\uparrow$0.1%) 58.7( $\downarrow$3.2%) 78.3( $\downarrow$0.0%) 77.7( $\downarrow$0.0%) 82.1( $\uparrow$0.0%) 0.6%±1.0%
Wiki. 84.1( $\uparrow$0.0%) 80.8( $\uparrow$0.2%) 88.8( $\downarrow$0.0%) 87.9( $\uparrow$0.0%) 75.2( $\uparrow$0.3%) 73.4( $\uparrow$0.5%) 89.5( $\downarrow$0.0%) 91.1( 0.0%) 83.3( $\uparrow$0.3%) 0.1%±0.2%
LastFM 76.4( $\downarrow$1.9%) 70.3( $\downarrow$2.5%) 78.5( $\downarrow$0.7%) 76.0( $\uparrow$0.0%) 72.2( $\downarrow$0.0%) 73.0( $\downarrow$0.2%) 72.5( $\uparrow$0.0%) 75.1( $\downarrow$0.0%) 82.1( $\uparrow$0.0%) 0.6%±1.0%
Myket 63.1( $\uparrow$0.4%) 61.7( $\downarrow$0.0%) 61.3( $\uparrow$0.0%) 56.4( $\downarrow$0.3%) 44.8( $\uparrow$0.0%) 51.1( $\uparrow$0.0%) 57.6( $\downarrow$0.1%) 58.7( $\downarrow$0.1%) 44.4( $\uparrow$0.0%) 0.1%±0.2%
Social 87.0( $\downarrow$1.0%) 90.0( $\downarrow$1.5%) 93.0( $\uparrow$0.0%) 94.9( $\uparrow$0.0%) 86.6( $\uparrow$0.2%) 80.8( $\uparrow$0.3%) 96.4( $\uparrow$0.1%) 95.5( $\uparrow$0.0%) 97.6( $\uparrow$0.1%) 0.4%±0.5%
Reddit 79.7( $\downarrow$0.1%) 78.9( 0.0%) 80.3( $\downarrow$0.1%) 78.3( $\downarrow$0.1%) 81.0( $\uparrow$0.1%) 73.4( $\downarrow$0.2%) 76.3( $\downarrow$0.0%) 77.2( $\downarrow$0.0%) 82.6( $\downarrow$0.0%) 0.1%±0.1%
$\mu\pm\sigma$ 1.5%±2.2% 2.9%±6.2% 0.3%±0.3% 0.1%±0.1% 0.1%±0.2% 1.1%±1.3% 0.1%±0.3% 0.2%±0.4% 0.2%±0.3%
Table 18: Average Precision scores as in Table 17 for discrete-time graphs. All results and standard deviations are listed in Table 20.
UN V. 53.3( $\downarrow$23.6%) 50.6( $\downarrow$28.2%) 49.9( $\downarrow$22.6%) 52.5( $\uparrow$1.7%) 52.6( $\uparrow$4.8%) 84.1( $\downarrow$0.4%) 52.4( $\downarrow$0.9%) 54.1( $\uparrow$1.4%) 60.0( $\uparrow$3.0%) 9.6%±11.6%
US L. 46.0( $\downarrow$4.2%) 62.5( $\downarrow$15.3%) 58.6( $\downarrow$28.5%) 71.0( $\uparrow$0.2%) 80.7( $\uparrow$0.0%) 63.2( $\downarrow$0.0%) 77.5( $\downarrow$0.1%) 86.5( $\uparrow$0.7%) 86.1( $\uparrow$0.6%) 5.5%±9.9%
UN Tr. 52.8( $\downarrow$10.0%) 49.6( $\downarrow$15.7%) 53.3( $\downarrow$9.6%) 59.1( $\uparrow$3.0%) 59.2( $\uparrow$3.3%) 79.0( $\downarrow$2.5%) 57.5( $\uparrow$3.5%) 65.8( $\uparrow$2.7%) 67.1( $\uparrow$3.3%) 6.0%±4.7%
Can. P. 52.3( $\uparrow$0.4%) 59.9( $\downarrow$6.4%) 69.7( $\downarrow$2.5%) 70.5( $\uparrow$0.1%) 66.6( $\uparrow$0.1%) 58.0( $\downarrow$2.8%) 67.0( $\uparrow$0.1%) 81.4( $\uparrow$0.2%) 82.2( $\downarrow$16.1%) 3.2%±5.3%
Flights 65.2( $\downarrow$2.3%) 63.4( $\downarrow$5.3%) 68.3( $\downarrow$0.8%) 73.5( $\downarrow$0.0%) 64.7( $\uparrow$0.7%) 70.3( 0.0%) 71.0( $\downarrow$0.1%) 71.9( $\downarrow$0.0%) 68.8( $\downarrow$0.6%) 1.1%±1.7%
Cont. 90.2( $\downarrow$2.2%) 95.1( $\downarrow$0.7%) 95.7( $\downarrow$0.7%) 96.0( $\downarrow$0.0%) 85.2( $\downarrow$0.0%) 88.7( $\downarrow$0.1%) 95.4( $\uparrow$0.4%) 93.5( $\downarrow$0.1%) 97.9( $\downarrow$0.0%) 0.5%±0.7%
$\mu\pm\sigma$ 7.1%±8.7% 11.9%±9.9% 10.8%±12.0% 0.8%±1.2% 1.5%±2.1% 1.0%±1.3% 0.8%±1.3% 0.8%±1.0% 3.9%±6.1%
Table 19: Test average precision scores for link forecasting and link prediction averaged over 5 runs with standard deviations on continuous-time temporal graphs.
Eval Dataset JODIE DyRep TGN TGAT CAWN EdgeBank TCL GraphMixer DyGFormer
Forec. Enron 69.8 ± 3.6 68.2 ± 4.0 65.3 ± 2.8 62.7 ± 0.8 66.8 ± 0.5 75.7 ± 0.0 70.7 ± 3.7 81.5 ± 0.6 77.1 ± 0.7
UCI 85.1 ± 1.8 56.6 ± 2.4 71.9 ± 0.9 69.0 ± 1.0 65.6 ± 0.7 66.9 ± 0.0 69.6 ± 0.8 85.9 ± 0.4 81.2 ± 0.9
MOOC 82.8 ± 5.1 76.2 ± 4.3 86.3 ± 2.7 84.4 ± 0.7 73.6 ± 0.9 58.7 ± 0.0 78.3 ± 0.6 77.7 ± 0.7 82.1 ± 9.8
Wiki. 84.1 ± 0.5 80.8 ± 0.3 88.8 ± 0.4 87.9 ± 0.2 75.2 ± 1.1 73.4 ± 0.0 89.5 ± 0.3 91.1 ± 0.2 83.3 ± 1.2
LastFM 76.4 ± 0.5 70.3 ± 2.0 78.5 ± 3.9 76.0 ± 0.7 72.2 ± 0.4 73.0 ± 0.0 72.5 ± 5.9 75.1 ± 1.2 82.1 ± 0.8
Myket 63.1 ± 1.8 61.7 ± 1.5 61.3 ± 2.1 56.4 ± 0.4 44.8 ± 0.3 51.1 ± 0.0 57.6 ± 2.3 58.7 ± 0.2 44.4 ± 1.7
Social 87.0 ± 6.1 90.0 ± 1.4 93.0 ± 2.4 94.9 ± 0.3 86.6 ± 0.1 80.8 ± 0.0 96.4 ± 0.2 95.5 ± 0.2 97.6 ± 0.1
Reddit 79.7 ± 0.4 78.9 ± 0.9 80.3 ± 0.6 78.3 ± 1.0 81.0 ± 0.5 73.4 ± 0.0 76.3 ± 0.6 77.2 ± 0.5 82.6 ± 0.8
Pred. Enron 69.0 ± 2.1 68.5 ± 4.3 65.0 ± 3.9 62.6 ± 0.6 66.6 ± 0.5 76.7 ± 0.0 70.1 ± 3.6 80.5 ± 0.6 76.5 ± 0.6
UCI 79.7 ± 3.0 48.0 ± 4.3 71.7 ± 1.1 69.1 ± 1.0 65.6 ± 0.7 64.9 ± 0.0 69.6 ± 0.8 86.0 ± 0.5 81.0 ± 1.0
MOOC 83.2 ± 4.5 76.6 ± 4.0 86.9 ± 2.2 84.3 ± 0.7 73.5 ± 1.0 60.6 ± 0.0 78.3 ± 0.5 77.7 ± 0.7 82.1 ± 9.8
Wiki. 84.1 ± 0.6 80.6 ± 0.3 88.8 ± 0.3 87.9 ± 0.2 75.0 ± 1.1 73.0 ± 0.0 89.5 ± 0.3 91.1 ± 0.2 83.1 ± 1.2
LastFM 77.9 ± 0.7 72.1 ± 1.9 79.1 ± 3.1 76.0 ± 0.6 72.2 ± 0.4 73.1 ± 0.0 72.4 ± 5.9 75.1 ± 1.2 82.1 ± 0.8
Myket 62.9 ± 1.8 61.7 ± 1.5 61.3 ± 2.1 56.6 ± 0.4 44.8 ± 0.3 51.1 ± 0.0 57.6 ± 2.2 58.7 ± 0.2 44.4 ± 1.7
Social 87.8 ± 5.3 91.3 ± 0.8 93.0 ± 2.6 94.9 ± 0.3 86.4 ± 0.1 80.5 ± 0.0 96.3 ± 0.2 95.4 ± 0.1 97.6 ± 0.1
Reddit 79.8 ± 0.4 78.9 ± 0.9 80.4 ± 0.6 78.4 ± 1.0 80.9 ± 0.5 73.6 ± 0.0 76.3 ± 0.6 77.2 ± 0.5 82.6 ± 0.8
Table 20: Test average precision scores for link forecasting and link prediction averaged over 5 runs with standard deviations on discrete-time temporal graphs.
Eval Dataset JODIE DyRep TGN TGAT CAWN EdgeBank TCL GraphMixer DyGFormer
Forec. UN V. 53.3 ± 1.2 50.6 ± 1.5 49.9 ± 4.4 52.5 ± 1.4 52.6 ± 1.9 84.1 ± 0.0 52.4 ± 0.9 54.1 ± 1.4 60.0 ± 2.0
UN Tr. 52.8 ± 3.1 49.6 ± 0.8 53.3 ± 1.7 59.1 ± 2.7 59.2 ± 1.7 79.0 ± 0.0 57.5 ± 1.9 65.8 ± 1.9 67.1 ± 2.7
Can. P. 52.3 ± 0.6 59.9 ± 6.5 69.7 ± 1.5 70.5 ± 1.8 66.6 ± 2.1 58.0 ± 0.0 67.0 ± 1.9 81.4 ± 0.5 82.2 ± 3.2
Flights 65.2 ± 2.6 63.4 ± 2.6 68.3 ± 2.2 73.5 ± 0.3 64.7 ± 0.8 70.3 ± 0.0 71.0 ± 0.4 71.9 ± 0.8 68.8 ± 2.0
Cont. 90.2 ± 5.2 95.1 ± 0.6 95.7 ± 1.0 96.0 ± 0.4 85.2 ± 0.2 88.7 ± 0.0 95.4 ± 0.6 93.5 ± 0.1 97.9 ± 0.1
Pred. UN V. 69.7 ± 1.5 70.5 ± 1.1 64.4 ± 6.5 51.6 ± 1.3 50.2 ± 1.4 84.5 ± 0.0 52.9 ± 1.4 53.3 ± 1.7 58.2 ± 1.5
US L. 48.0 ± 1.0 73.8 ± 2.4 81.9 ± 2.3 70.9 ± 8.7 80.7 ± 3.6 63.2 ± 0.0 77.6 ± 4.4 85.8 ± 2.0 85.6 ± 1.2
UN Tr. 58.7 ± 3.7 58.9 ± 1.7 58.9 ± 1.5 57.4 ± 2.6 57.4 ± 2.3 81.0 ± 0.0 55.6 ± 1.5 64.1 ± 1.7 64.9 ± 2.6
Can. P. 52.1 ± 0.4 64.0 ± 1.7 71.5 ± 1.8 70.4 ± 1.7 66.5 ± 2.4 59.7 ± 0.0 67.0 ± 1.9 81.2 ± 0.4 98.0 ± 0.5
Flights 66.7 ± 3.6 66.9 ± 1.9 68.9 ± 2.0 73.5 ± 0.3 64.2 ± 1.0 70.3 ± 0.0 71.0 ± 0.4 71.9 ± 0.8 69.3 ± 2.0
[…]

### v2

@@ G.3 Average precision
Can. P. 52.8 ± 0.5 61.8 ± 1.1 68.5 ± 2.1 67.7 ± 1.7 63.9 ± 1.3 63.8 ± 0.0 64.2 ± 2.0 77.2 ± 0.4 97.1 ± 0.7
Flights 66.7 ± 3.3 66.8 ± 1.6 68.3 ± 1.8 72.7 ± 0.2 65.0 ± 1.3 70.5 ± 0.0 70.8 ± 0.5 71.2 ± 0.7 68.6 ± 1.7
Cont. 94.2 ± 1.3 95.5 ± 0.3 96.3 ± 1.1 96.0 ± 0.3 84.5 ± 0.2 88.8 ± 0.0 95.0 ± 0.6 94.2 ± 0.1 97.7 ± 0.1
Avg. Rank 6.5 5.3 4.0 5.0 7.5 4.8 5.7 3.5 2.7
G.4 AUC-ROC Scores over Time
We present the AUC-ROC scores across the test time for all models and continuous-time datasets with an average relative change in performance (rightmost column in Table 3) larger than one, i.e. Enron (Figure 9), UCI (Figure 10), MOOC (Figure 11), and LastFM (Figure 12). We observe that the patterns seen in Figure 4 for the Enron and MOOC dataset extend to all models. For UCI in Figure 10, we can see the AUC-ROC scores for many time windows and comparably fewer batches. This is because the test set for this dataset contains 15% of all temporal edges but two thirds of the whole observation time.
Figure 12 showing the LastFM dataset contains several outliers in the first half of the test time where the performance is better compared to the average. These outliers correspond to the peaks in activity shown in Figure 5. Similar to what can be observed for the one anomaly in the Enron dataset, these peaks in activity lead to many batches but only a small number of time windows. This is because these bursts occur in a short period of time, leading to an overrepresentation of these outliers using the batch-based approach. At the end of the LastFM dataset, fewer edges occur leading to a less reliable model performance which is underrepresented in the batch-based evaluation due to the fewer number of edges.
Figure 9: Model performance of Enron dataset over time which is visualised by the AUC-ROC scores for each batch and time window individually. The time (x-axis) is limited to the test set.
Figure 10: Model performance of UCI over time.
Figure 11: Model performance of MOOC over time.
Figure 12: Model performance of LastFM over time.
G.5 Investigating Information Leakage
The problem of information leakage can give memory-based models an unfair advantage during evaluation because these models obtain more information than non-memory-based models due to this leakage. However, whether this leaked information is useful for making better predictions depends on the data. We assume that the order of the edges within each snapshot plays an important role in many cases where the leaked information is useful. The data could, for example, retain an implicit temporal ordering if the data was obtained by coarse-graining a continuous-time graph into snapshots. The edges inside the snapshots could also be sorted by some ID, like the node ID of the source node. This would mean that all occurrences of each node as a source node would be consecutive, making it very unlikely that each node appears again as a source node after it was already seen in a previous batch of the same snapshot.
In Figure 13, we show the AUC-ROC scores of each batch individually in the evaluated order and highlight the borders of each snapshot with grey dashed lines. We observe upwards trends in the red regression lines for the US Legislators (DyRep, TGN), the UN Trade (all models), and the Flights (JODIE, DyRep) datasets. This suggests that these datasets contain edges in an order that simplifies the prediction for the indicated models given knowledge from previous batches. Thus, shuffling the data within each snapshot should destroy the useful leaked information and lead to a drop in performance. We report the changes in performance between (i) the models that are evaluated based on edges in the order found in the dataset and (ii) the models evaluated on the datasets with the edges inside each snapshot shuffled in Table 18.
The tables contain the relative change in performance for both evaluation approaches: the batch-based link prediction and our window-based link forecasting. We observe that the performance of the batch-based approach drops for the above-identified dataset and model combinations, suggesting that the edges are indeed ordered in such a way that the earlier batches in a snapshot provide information about later batches, making the prediction easier. Our time-window-based approach prevents this information leakage. Thus, the performance after shuffling is similar to the performance without shuffling because no information, whether useful or not, is leaked.
Table 18: Relative change of the performance between test sets with the order of edges within snapshots as provided by the discrete-time datasets and test sets with edges shuffled within each snapshot. We compare the evaluation approaches of our window-based link forecasting with batch-based link prediction on all memory-based models.
(a) AUC-ROC scores
Model JODIE DyRep TGN
Eval Forec. Pred. Forec. Pred. Forec. Pred.
UN V. $\downarrow$1.6$\downarrow$20.0$\downarrow$7.1$\downarrow$21.4$\downarrow$1.7$\downarrow$15.3%
US L. $\downarrow$0.1$\downarrow$3.5$\downarrow$0.1$\downarrow$10.9$\downarrow$0.6$\downarrow$10.4%
UN Tr. $\downarrow$0.2$\downarrow$13.8$\uparrow$0.1$\downarrow$15.5$\downarrow$3.5$\downarrow$14.2%
Can. P. $\downarrow$0.3$\uparrow$4.3$\downarrow$4.4$\downarrow$0.1$\downarrow$14.0$\downarrow$10.1%
Flights $\uparrow$1.8$\downarrow$1.4$\uparrow$1.3$\downarrow$2.0$\uparrow$0.3$\uparrow$0.1%
Cont. $\downarrow$0.9$\downarrow$1.2$\uparrow$0.1$\downarrow$1.1$\downarrow$0.1$\downarrow$0.2%
$\mu$ 0.8% 7.4% 2.2% 8.5% 3.4% 8.4%
(b) Average precision scores
Model JODIE DyRep TGN
Eval Forec. Pred. Forec. Pred. Forec. Pred.
UN V. $\downarrow$4.2$\downarrow$18.4$\downarrow$6.1$\downarrow$20.5$\downarrow$1.5$\downarrow$13.1%
US L. $\downarrow$0.0$\downarrow$2.2$\downarrow$0.1$\downarrow$11.0$\downarrow$1.5$\downarrow$10.5%
UN Tr. $\downarrow$0.2$\downarrow$10.0$\uparrow$0.9$\downarrow$11.4$\downarrow$3.3$\downarrow$9.9%
Can. P. $\downarrow$0.2$\uparrow$3.1$\downarrow$5.3$\downarrow$0.1$\downarrow$17.6$\downarrow$9.1%
Flights $\uparrow$1.0$\downarrow$2.0$\uparrow$0.9$\downarrow$3.7$\uparrow$0.4$\uparrow$0.7%
Cont. $\downarrow$2.3$\downarrow$3.1$\downarrow$0.2$\downarrow$2.4$\downarrow$0.2$\downarrow$0.6%
$\mu$ 1.3% 6.5% 2.2% 8.2% 4.1% 7.3%
(a) UN Vote
(b) US Legislators
(c) UN Trade
(d) Canadian Parliament
(e) Flights
[…]